OpenStudy (anonymous):

Solve for the indicated variable. A=h/2 (a+b); for h

OpenStudy (anonymous):

2A / h = a+b subtract b a=(2A / h) - b

OpenStudy (anonymous):

@xxccjxx that is incorrect.

OpenStudy (noorik.):

is h/[2(a+b)] or (h/2)(a+b)?

OpenStudy (anonymous):

no, that's incorrect too

OpenStudy (noorik.):

i'm asking you about the form of the question, try typing it using the "Equation", maybe it'll help us to understand the question correclty.

OpenStudy (anonymous):



OpenStudy (noorik.):

then \[h=\frac{ 2A }{ a+b }\]

OpenStudy (anonymous):

thanks so much @NooriK. can you help me with 2 more?

OpenStudy (noorik.):

you are welcome, yes i can help you, what are the problems?

OpenStudy (anonymous):

\[C= 2\pi r \] ; for r

OpenStudy (noorik.):

\[r = \frac{ C }{ 2\pi }\] \[{\pi} = 3.14\]

OpenStudy (anonymous):

P = 2L + 2W; for L

OpenStudy (noorik.):

P = 2L + 2W P - 2W = 2L L = (P-2W)/2

OpenStudy (anonymous):

thanks so much :)